# Nova: Generative Language Models for Assembly Code with Hierarchical Attention and Contrastive Learning

### v6

@@ 1 Introduction
Binary code plays an irreplaceable role in the security domain, being the foundation of crucial tasks including vulnerability detection (Güler et al., 2019; Duan et al., 2020; Chen et al., 2022b), malware detection (Spensky et al., 2016; Aonzo et al., 2023; Xu et al., 2014), binary recovery (Su et al., 2024; Zhang et al., 2021; Chen et al., 2022c), and legacy software maintenance (Carbone et al., 2009; Carlini et al., 2015; Martin et al., 2010). For example, when performing tasks such as identifying attacks and malware, security analysts often only have access to assembly, i.e., the human-readable representation of binary code, which is extremely difficult to understand (Su et al., 2024; Zhang et al., 2021; Chen et al., 2022c). Thus, combined with the increasing sophistication of cybercrime that poses significant threats worldwide (e.g., cybercrime is predicted to cost the world \$10.5 trillion annually by 2025 (Sausalito, 2020)), effective binary analysis techniques are in high demand.
@@
Large language models pre-trained on source code have brought improvement in various software development domains (Chen et al., 2022a; Liu et al., 2023a; Chen et al., 2023; Le et al., 2022; Jiang et al., 2023; Xia et al., 2023). However, these LLMs are not designed for or trained with assembly corpus, not achieving their full potential on binary code analysis tasks such as binary code similarity (Wang et al., 2022; Xu et al., 2023a), malware detection (Su et al., 2024), and binary code decompilation (Tan et al., 2024; Armengol-Estapé et al., 2024; Hosseini & Dolan-Gavitt, 2022).
Existing work applying LLMs on assembly code mainly piggybacks on encoder-style LLMs (Wang et al., 2022; Su et al., 2024; Xu et al., 2023a), unable to benefit from the more extensive pre-training, updated architectures, scaling of state-of-the-art generative LLMs. Other work using generative LLMs for decompilation shows a low unit test passing rate of the decompiled programs (Tan et al., 2024; Armengol-Estapé et al., 2024).
@@ 2.1 Binary Models
Machine learning models are widely used in binary program analysis tasks. However, these models are typically designed for specific tasks such as binary code similarity detection (Pei et al., 2020; Xu et al., 2023a; Wang et al., 2022; Xu et al., 2017; Ding et al., 2019), variable name prediction (Chen et al., 2022c; Xu et al., 2023b; Zhang et al., 2021; He et al., 2018; Lacomis et al., 2020), binary code type inference (Pei et al., 2021), and so on (Chen et al., 2022d; Liu et al., 2023b; Hosseini & Dolan-Gavitt, 2022).
Recent techniques have started to pre-train foundation LLMs for binaries. CodeArt (Su et al., 2024) pre-trains encoder-style LLMs with a regularized attention design to better encode assembly code semantics. SLaDe (Armengol-Estapé et al., 2024) trains BART (Lewis et al., 2019) models on assembly. Meta LLMCompiler (Cummins et al., 2024) train CodeLlama models on LLVM IR to optimize binary code. LLM4Decompile (Tan et al., 2024) trains DeepSeekCoder with assembly for binary code decompilation. However, CodeArt does not generalize to generation tasks due to its encoder architecture. LLMCompiler trained on LLVM IR cannot be effectively transferred to assembly code. SLaDe and LLM4Decompile are limited in performance due to a lack of special designs for assembly. In contrast, Nova proposes hierarchical attention and contrastive learning objectives, outperforming existing techniques on both understanding (binary code similarity detection) and generation (binary code decompilation) tasks.
@@ 2.2 Large Source-Code Models
LLMs demonstrate promising results on many code-related tasks, such as code generation (Chen et al., 2022a; Liu et al., 2023a; Chen et al., 2023; Le et al., 2022; Yue et al., 2021; Chen et al., 2021; Nijkamp et al., 2022; Fried et al., 2023; Rozière et al., 2023; Guo et al., 2024; Lozhkov et al., 2024; Hui et al., 2024), bug fixing (Jiang et al., 2023; Xia et al., 2023) and vulnerability fixing (Wu et al., 2023; Steenhoek et al., 2023; He & Vechev, 2023). The advances in using LLMs are attributed to the knowledge learned from massive source code and natural language text in their training datasets (Touvron et al., 2023; OpenAI, 2023). Nova is designed and trained for assembly, which has unique challenges such as low information density and diverse optimization.
@@ 2.3 Attention Mechanism
Standard self-attention is widely used in transformer architecture (Vaswani et al., 2017) to capture soft dependencies between tokens in the input. Many special attention mechanisms have been designed for better learning in various scenarios (Yang et al., 2016; Huang et al., 2024). LongCoder (Guo et al., 2023) combines window attention and global attention (attention sink (Xiao et al., 2024)) to handle long input of source code. We have shown that LongCoder’s window attention is less effective than Nova’s on assembly code. CAST (Shi et al., 2021) is a new neural architecture that splits the abstract syntax tree (AST) of source code into subtrees, encodes the subtrees, and aggregates to the final encoding. PA-former (Chai & Li, 2024) is a new neural architecture that constructs source code as pyramid input based on their AST structure and contains a pyramid attention mechanism to calculate the features in a hierarchical way. HierarchyNet (Nguyen et al., 2023) is a neural architecture that considers source code AST, data flow, and control flow graphs. Similarly, it cannot be applied to assembly code. Different from CAST, PA-former, and HierarchiyNet, Nova’s attention design is for assembly code, is more lightweight and can be plugged into any pre-trained generative LLM.
@@ 3 Approach
Figure 2 presents the overall approach of Nova. We build Nova on top of foundation models for source code (Guo et al., 2024) to utilize their source code and natural language generation ability. We first collect large assembly corpora (Section 3.1). Section 3.2 describes Nova’s hierarchical attention design. With the collected assembly corpora, we then pretrain Nova with language modeling and contrastive learning objectives (Section 3.3). Then, we fine-tune Nova on two important downstream tasks, binary code decompilation, and binary code similarity detection (Sections 3.4 and 3.5), to prove Nova’s effectiveness and benefits to the binary research domain.
@@ 3.1 Data Collection
We derive our X86-64 assembly functions dataset from two source code corpora: C functions in The-Stack (Li et al., 2023) and AnghaBench (da Silva et al., 2021). We compile the C programs into executables using gcc with different optimization levels (i.e., O0, O1, O2 and O3), strip the executables to remove debug information, and disassemble them into X86-64 assembly code using objdump. We treat every function as a separate data point. Table 3 shows the number of C functions in the two original datasets, and the number of X86-64 assembly functions we collected from them.
@@ 3.2 Hierarchical Self-Attention
Nova uses hierarchical self-attention that is specially designed to learn the low-information-density semantics in the long sequence of assembly code. Specifically, Nova learns the assembly code in an hierarchical way by providing a modified attention mask. Different from standard token-level attentions (Vaswani et al., 2017; Radford & Narasimhan, 2018; Radford et al., 2019; Brown et al., 2020), our hierarchical self-attention contains three categories at different levels of granularity.
@@ 3.3 Contrastive Learning
The syntax gap between assembly code and source code, and syntax diversity between differently-optimized assembly code make LLMs struggle to distinguish the semantics behind the syntax. Nova adopts contrastive learning technique (Gao et al., 2021) during pre-training to train LLMs to encode assembly code meaningfully w.r.t semantics. The standard pre-training objective is language modeling by minimizing the negative likelihood of code in the pre-training corpus (Radford & Narasimhan, 2018), notated as $L_{lm}$. In addition, Nova is pre-trained with two new objectives, $L_{fcl}$ for functionality contrastive learning and $L_{ocl}$ for optimization contrastive learning.
@@
Let $e^{s}_{f}$ be the embedding of function $f$ in $s$ form ($s=-1$ for source code, and $s\in[0,1,2,3]$ for O0 to O3 optimized assembly). For simplicity, let $S=[-1,0,1,2,3]$ be the domain of $s$. Functionality CL optimizes Nova’s embeddings to satisfy the constraint:
[…]
@@ 3.4 Task 1: Binary Code Decompilation
[…]
@@ 3.5 Task 2: Binary Code Similarity Detection
Binary code similarity detection (BCSD) aims to measure the similarity between two binary code snippets (Wang et al., 2022; Su et al., 2024), which is the foundation of various applications such as plagiarism detection (Luo et al., 2014; Sæbjørnsen et al., 2009) and vulnerability detection (David & Yahav, 2014; David et al., 2018; 2017; 2016).
@@
[…]
We follow previous work (Su et al., 2024) to let $s$ be O0-assembly and $t$ be O3-assembly, which is the hardest setting.
@@ 4.1 Pre-Training
We use the C and X86-64 assembly functions collected from AnghaBench and The-Stack for pre-training. We pre-train Nova starting from DeepSeek-Coder (Guo et al., 2024), and the hierarchical attention is applied on half of the attention heads to balance between its effectiveness and the existing knowledge in the standard attention layers (Justified in Appendix A.3). Nova is pre-trained with language modeling for one epoch, followed by contrastive learning objectives for another epoch.
@@ 4.2 Fine-Tuning for Binary Code Decompilation
Test Data: We use HumanEval-Decompile (Tan et al., 2024) as the test benchmark, which was not used in training. HumanEval-Decompile is derived from the C language adaptation of the HumanEval (Chen et al., 2021) benchmark and contains 164 C functions, each compiled with O0O3 optimization flags and disassembled into X86-64 assembly.
Baselines: Nova is compared with existing SOTA LLM4Decompile (Tan et al., 2024), Meta LLMCompiler (Cummins et al., 2024), other open-sourced general code LLMs (Rozière et al., 2023; Lozhkov et al., 2024; Guo et al., 2024; Hui et al., 2024; Mishra et al., 2024; Guo et al., 2023), and commercial LLMs GPT-3.5-Turbo and GPT-4o). LLM4Decompile trains DeepSeekCoder using the same AnghaBench corpus for binary decompilation. Meta LLMCompiler trains CodeLlama models using LLVM IRs, X86, and ARM assembly code to optimize and translate binary code.
Evaluation: We let each model sample 20 decompilations per assembly function, using the temperature of 0.2 and top_p of 0.95 (Chen et al., 2021). Except for LLM4Decompile and Nova that are fine-tuned for binary code decompilation, we provide all other baselines with three-shot examples for few-shot learning (Brown et al., 2020). The decompilations are executed with the test cases to verify the functional correctness. Finally, Pass@1 and Pass@10 (Chen et al., 2021) are reported.
@@ 4.3 Fine-Tuning for Binary Code Similarity Detection
Training Data: To compare Nova with existing works on BCSD fairly (Wang et al., 2022; Su et al., 2024), we use the same dataset, BinaryCorp-3M (Wang et al., 2022), as the fine-tuning data for BCSD, which contains the O0 and O3 assembly of 224,606 functions.
Test Data: Following existing work (Su et al., 2024; Xu et al., 2023a), we use real-world benchmarks, Binutils, Curl, ImageMagick, SQLite, OpenSSL, and Putty, as the test benchmarks, which are nonexistent in the training data.
Baselines: Nova is compared with jTrans (Wang et al., 2022), DiEmph (Xu et al., 2023a) and CodeArt (Su et al., 2024). jTrans is a Transformer (Vaswani et al., 2017) encoder trained on binaries with masked token prediction and jump target prediction tasks. DiEmph uses an instruction deemphasis technique to prevent the model from learning instruction distribution biases introduced by compilers. CodeArt proposes a regularized attention mask for encoder models to capture instructional semantics and data dependencies.
@@ ACKNOWLEDGMENTS
This research was supported in part by NSF 1901242 and 2006688, a CFI fund, and J.P. Morgan AI Faculty Research Awards. Any opinions, findings, and conclusions in this paper are those of the authors only and do not necessarily reflect the views of our sponsors.
@@ A.2 Experimental Setup Details
For the evaluation of BCSD, we reuse the framework provided by CodeArt Su et al. (2024) to evaluate the binary code similarity detection results once Nova produces the embeddings for functions in the test dataset. For each one of the $K$ functions, the O0 assembly function is used as the query to calculate the cosine similarity between its embedding and the embeddings of all the $K$ O3 assembly functions. Note that Nova does not normalize the embeddings of assembly functions during training and uses $l_{2}$ distance to calculate the $L_{BCSD}$, which optimize Nova to embed the O0 and O3 assembly functions from the same source code have the smallest $l_{2}$ distances. When using CodeArt’s framework for evaluation which ranks the similarity using the cosine similarity, we normalize Nova’s embedding for each assembly function since $l_{2}$ distance after normalization keeps the same order as cosine similarity (smaller $l_{2}$ distance means higher cosine similarity).

### v7

@@ 1 Introduction
Binary code plays an irreplaceable role in the security domain, being the foundation of crucial tasks including vulnerability detection (vul-motivate-1; vul-motivate-2; vul-motivate-3), malware detection (malware-motivate-1; malware-motivate-2; malware-motivate-3), binary recovery (codeart; osprey; dirty), and legacy software maintenance (legacy-motivate-1; legacy-motivate-2; legacy-motivate-3). For example, when performing tasks such as identifying attacks and malware, security analysts often only have access to assembly, i.e., the human-readable representation of binary code, which is extremely difficult to understand (codeart; osprey; dirty). Thus, combined with the increasing sophistication of cybercrime that poses significant threats worldwide (e.g., cybercrime is predicted to cost the world \$10.5 trillion annually by 2025 (cyber-attack)), effective binary analysis techniques are in high demand.
@@
Large language models pre-trained on source code have brought improvement in various software development domains (related-generation-1; related-generation-2; related-generation-3; related-generation-4; related-apr-1; related-apr-2). However, these LLMs are not designed for or trained with assembly corpus, not achieving their full potential on binary code analysis tasks such as binary code similarity (jtrans; xu2023diemph), malware detection (codeart), and binary code decompilation (llm4decompile; slade; related-decompile-1).
Existing work applying LLMs on assembly code mainly piggybacks on encoder-style LLMs (jtrans; codeart; xu2023diemph), unable to benefit from the more extensive pre-training, updated architectures, scaling of state-of-the-art generative LLMs. Other work using generative LLMs for decompilation shows a low unit test passing rate of the decompiled programs (llm4decompile; slade).
@@ 2.1 Binary Models
Machine learning models are widely used in binary program analysis tasks. However, these models are typically designed for specific tasks such as binary code similarity detection (pei2020trex; xu2023diemph; jtrans; related-bcsd-1; asm2vec), variable name prediction (dirty; xu2023lmpa; osprey; debin; dire), binary code type inference (pei2021stateformer), and so on (related-reverse-1; related-decompile-2; related-decompile-1).
Recent techniques have started to pre-train foundation LLMs for binaries. CodeArt (codeart) pre-trains encoder-style LLMs with a regularized attention design to better encode assembly code semantics. SLaDe (slade) trains BART (bart) models on assembly. Meta LLMCompiler (llm-compiler) train CodeLlama models on LLVM IR to optimize binary code. LLM4Decompile (llm4decompile) trains DeepSeekCoder with assembly for binary code decompilation. However, CodeArt does not generalize to generation tasks due to its encoder architecture. LLMCompiler trained on LLVM IR cannot be effectively transferred to assembly code. SLaDe and LLM4Decompile are limited in performance due to a lack of special designs for assembly. In contrast, Nova proposes hierarchical attention and contrastive learning objectives, outperforming existing techniques on both understanding (binary code similarity detection) and generation (binary code decompilation) tasks.
@@ 2.2 Large Source-Code Models
LLMs demonstrate promising results on many code-related tasks, such as code generation (related-generation-1; related-generation-2; related-generation-3; related-generation-4; codet5; codex; codegen; incoder; codellama; deepseekcoder; starcoder2; qwencoder), bug fixing (related-apr-1; related-apr-2) and vulnerability fixing (related-vul-1; related-vul-2; related-vul-4). The advances in using LLMs are attributed to the knowledge learned from massive source code and natural language text in their training datasets (llama; openai2023gpt4). Nova is designed and trained for assembly, which has unique challenges such as low information density and diverse optimization.
@@ 2.3 Attention Mechanism
Standard self-attention is widely used in transformer architecture (attention) to capture soft dependencies between tokens in the input. Many special attention mechanisms have been designed for better learning in various scenarios (hierarchical-attention; long-context-survey). LongCoder (longcoder) combines window attention and global attention (attention sink (streamingllm)) to handle long input of source code. We have shown that LongCoder’s window attention is less effective than Nova’s on assembly code. CAST (cast) is a new neural architecture that splits the abstract syntax tree (AST) of source code into subtrees, encodes the subtrees, and aggregates to the final encoding. PA-former (pyramid-attention) is a new neural architecture that constructs source code as pyramid input based on their AST structure and contains a pyramid attention mechanism to calculate the features in a hierarchical way. HierarchyNet (hierarchy-net) is a neural architecture that considers source code AST, data flow, and control flow graphs. Similarly, it cannot be applied to assembly code. Different from CAST, PA-former, and HierarchiyNet, Nova’s attention design is for assembly code, is more lightweight and can be plugged into any pre-trained generative LLM.
@@ 3 Approach
Figure 2 presents the overall approach of Nova. We build Nova on top of foundation models for source code (deepseekcoder) to utilize their source code and natural language generation ability. We first collect large assembly corpora (Section 3.1). Section 3.2 describes Nova’s hierarchical attention design. With the collected assembly corpora, we then pretrain Nova with language modeling and contrastive learning objectives (Section 3.3). Then, we fine-tune Nova on two important downstream tasks, binary code decompilation, and binary code similarity detection (Sections 3.4 and 3.5), to prove Nova’s effectiveness and benefits to the binary research domain.
@@ 3.1 Data Collection
We derive our X86-64 assembly functions dataset from two source code corpora: C functions in The-Stack (starcoder) and AnghaBench (anghabench). We compile the C programs into executables using gcc with different optimization levels (i.e., O0, O1, O2 and O3), strip the executables to remove debug information, and disassemble them into X86-64 assembly code using objdump. We treat every function as a separate data point. Table 3 shows the number of C functions in the two original datasets, and the number of X86-64 assembly functions we collected from them.
@@ 3.2 Hierarchical Self-Attention
Nova uses hierarchical self-attention that is specially designed to learn the low-information-density semantics in the long sequence of assembly code. Specifically, Nova learns the assembly code in an hierarchical way by providing a modified attention mask. Different from standard token-level attentions (attention; gpt-1; gpt-2; gpt-3), our hierarchical self-attention contains three categories at different levels of granularity.
@@ 3.3 Contrastive Learning
The syntax gap between assembly code and source code, and syntax diversity between differently-optimized assembly code make LLMs struggle to distinguish the semantics behind the syntax. Nova adopts contrastive learning technique (simcse) during pre-training to train LLMs to encode assembly code meaningfully w.r.t semantics. The standard pre-training objective is language modeling by minimizing the negative likelihood of code in the pre-training corpus (gpt-1), notated as $L_{lm}$. In addition, Nova is pre-trained with two new objectives, $L_{fcl}$ for functionality contrastive learning and $L_{ocl}$ for optimization contrastive learning.
@@
Let $e^{s}_{f}$ be the embedding of function $f$ in $s$ form (s=1s=-1 for source code, and $s\in[0,1,2,3]$ for O0 to O3 optimized assembly). For simplicity, let S=[1,0,1,2,3]S=[-1,0,1,2,3] be the domain of $s$. Functionality CL optimizes Nova’s embeddings to satisfy the constraint:
$\forall f_{i}\in F,\;\mathop{max}\limits_{s,t\in S}(\mathop{d}(e^{s}_{f_{i}},e^{t}_{f_{i}}))<\mathop{min}\limits_{\begin{subarray}{c}s,t\in S\\ f_{j}\neq f_{i}\in F\end{subarray}}(\mathop{d}(e^{s}_{f_{i}},e^{t}_{f_{j}}))$
, where $\mathop{d}$ calculates the $l_{2}$ distance between two embeddings and $F$ is the full set of functions in the training corpus.
@@ 3.4 Task 1: Binary Code Decompilation
Binary code decompilation (BCD) helps developers to understand binary code by recovering binary code into more readable high-level source code (e.g., C programs) (coda; neutron; slade; llm4decompile). The input to the model for BCD is formatted as an instruction prompt (notated by $\mathbf{p}$): # This is the assembly code with {opt} optimization: $\mathbf{asm}$}, where “opt” is the optimization-level applied to the assembly and $\mathbf{asm}$” is the assembly code to decompile. Nova is fine-tuned to generate the expected source code function $\mathbf{src}$ following the instruction prompt. The fine-tuning objective is minimizing the loss: $L_{bcd}=-\mathop{log}\mathop{P}(\mathbf{src}|\mathbf{p})$.
@@ 3.5 Task 2: Binary Code Similarity Detection
Binary code similarity detection (BCSD) aims to measure the similarity between two binary code snippets (jtrans; codeart), which is the foundation of various applications such as plagiarism detection (plagiarism-1; plagiarism-2) and vulnerability detection (binary-vul-1; binary-vul-2; binary-vul-3; binary-vul-4).
@@
$L_{BCSD}=-\mathop{log}\sum\limits_{1\leq i\leq K}^{f^{q}:=f^{p}_{i}}\left(1-\frac{\mathop{exp}\left(\mathop{d}(e^{s}_{f^{q}},e^{t}_{f^{p}_{i}})\right)}{\sum_{1\leq j\leq K}\mathop{exp}\Big(\mathop{d}(e^{s}_{f^{q}},e^{t}_{f^{p}_{j}})\Big)}\right)$
We follow previous work (codeart) to let $s$ be O0-assembly and $t$ be O3-assembly, which is the hardest setting.
@@ 4.1 Pre-Training
We use the C and X86-64 assembly functions collected from AnghaBench and The-Stack for pre-training. We pre-train Nova starting from DeepSeek-Coder (deepseekcoder), and the hierarchical attention is applied on half of the attention heads to balance between its effectiveness and the existing knowledge in the standard attention layers (Justified in Appendix A.3). Nova is pre-trained with language modeling for one epoch, followed by contrastive learning objectives for another epoch.
@@ 4.2 Fine-Tuning for Binary Code Decompilation
Test Data: We use HumanEval-Decompile (llm4decompile) as the test benchmark, which was not used in training. HumanEval-Decompile is derived from the C language adaptation of the HumanEval (codex) benchmark and contains 164 C functions, each compiled with O0O3 optimization flags and disassembled into X86-64 assembly.
Baselines: Nova is compared with existing SOTA LLM4Decompile (llm4decompile), Meta LLMCompiler (llm-compiler), other open-sourced general code LLMs (codellama; starcoder2; deepseekcoder; qwencoder; granite; longcoder), and commercial LLMs GPT-3.5-Turbo and GPT-4o). LLM4Decompile trains DeepSeekCoder using the same AnghaBench corpus for binary decompilation. Meta LLMCompiler trains CodeLlama models using LLVM IRs, X86, and ARM assembly code to optimize and translate binary code.
Evaluation: We let each model sample 20 decompilations per assembly function, using the temperature of 0.2 and top_p of 0.95 (codex). Except for LLM4Decompile and Nova that are fine-tuned for binary code decompilation, we provide all other baselines with three-shot examples for few-shot learning (gpt-3). The decompilations are executed with the test cases to verify the functional correctness. Finally, Pass@1 and Pass@10 (codex) are reported.
@@ 4.3 Fine-Tuning for Binary Code Similarity Detection
Training Data: To compare Nova with existing works on BCSD fairly (jtrans; codeart), we use the same dataset, BinaryCorp-3M (jtrans), as the fine-tuning data for BCSD, which contains the O0 and O3 assembly of 224,606 functions.
Test Data: Following existing work (codeart; xu2023diemph), we use real-world benchmarks, Binutils, Curl, ImageMagick, SQLite, OpenSSL, and Putty, as the test benchmarks, which are nonexistent in the training data.
Baselines: Nova is compared with jTrans (jtrans), DiEmph (xu2023diemph) and CodeArt (codeart). jTrans is a Transformer (attention) encoder trained on binaries with masked token prediction and jump target prediction tasks. DiEmph uses an instruction deemphasis technique to prevent the model from learning instruction distribution biases introduced by compilers. CodeArt proposes a regularized attention mask for encoder models to capture instructional semantics and data dependencies.
@@ ACKNOWLEDGMENTS
This research was supported in part by NSF 1901242 and 2006688, a CFI fund, and J.P. Morgan AI Faculty Research Awards. This work also used Anvil at Purdue University through allocation CIS240304 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program (anvil), which is supported by NSF grants 2138259, 2138286, 2138307, 2137603, and 2138296. Any opinions, findings, and conclusions in this paper are those of the authors only and do not necessarily reflect the views of our sponsors.
@@ A.2 Experimental Setup Details
For the evaluation of BCSD, we reuse the framework provided by CodeArt codeart to evaluate the binary code similarity detection results once Nova produces the embeddings for functions in the test dataset. For each one of the $K$ functions, the O0 assembly function is used as the query to calculate the cosine similarity between its embedding and the embeddings of all the $K$ O3 assembly functions. Note that Nova does not normalize the embeddings of assembly functions during training and uses $l_{2}$ distance to calculate the $L_{BCSD}$, which optimize Nova to embed the O0 and O3 assembly functions from the same source code have the smallest $l_{2}$ distances. When using CodeArt’s framework for evaluation which ranks the similarity using the cosine similarity, we normalize Nova’s embedding for each assembly function since $l_{2}$ distance after normalization keeps the same order as cosine similarity (smaller $l_{2}$ distance means higher cosine similarity).